OpenStudy (anonymous):

Which type of triangle is formed with the points A(1, 7), B(-2, 2), and C(4, 2) as its vertices?

OpenStudy (wwe123):

A(-1,2,3), B(-1,3,5), C(0,-1,1) AB=OB-OA AB=(-1,3,5) - (-1,2,3) AB= (0,1,2) AC=OC-0A AC=(0,-1,1) - (-1,2,3) AC=(1,-3,-2) Using the scalar product cosx= (AB. AC)/(|AB| |AC|) cosx=(0,1,2) . (1,-3,2)/ sqrt(0^2 +1^2 +2^2) sqrt(1^2+3^2+2^2) x=83.1 degrees |A| = sqrt14 |B| = sqrt35 A=1/2 ab sinC A= 1/2(sqrt14sqrt35)sin83.1 A= 10.99 units^2